OpenStudy (anonymous):

Use an addition or subtraction formula to find the exact value of the expression. sin(285)

OpenStudy (anonymous):

Well, lets assume you know the 'obvious' values of sin - {0, 30, 45, 60, 90} + 360n. How can you use them to make 285?

OpenStudy (anonymous):

add them?

OpenStudy (anonymous):

Indeed. Use sin(a+b) with a couple of them and you should be able to work it out. I guess the main problem is how to do it most efficiently.

OpenStudy (anonymous):

so what do i add?

OpenStudy (anonymous):

That's where i'm cofused

OpenStudy (anonymous):

Something to consider is that sin (180 - x) = sin x (the best way is probably to add 2 values of sin between 100 and 180

OpenStudy (anonymous):

¬_¬ Sin 150 = Sin 30 Sin 135 = Sin 45

OpenStudy (anonymous):

I do not get it.

OpenStudy (anonymous):

Do you not get why that is true?

OpenStudy (anonymous):

Or do you not get that sin(150 + 135) = sin 285

OpenStudy (anonymous):

both

OpenStudy (anonymous):

Well, if 0 < x < 180 , then sin (180 - x) = sin x Hence the above claim. And 135 + 150 = 285 (clearly) sin(150 + 135) = sin(150)cos(135) + sin(135)cos(150) = sin(30)cos(135) + sin(45)(cos(150) All you need to do is use a SIMILAR rule one the cos values and you're done. I'll leave that to you. NB this is how I'd do it, but there may be a quicker way.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Gah, actually I prefer: Sin 285 = sin (285 - 360) = sin(-75) Sin is an odd function => sin(-75) = -sin(75) sin 75 = sin(30+45) = blah. -sin 35 = -blah.

OpenStudy (anonymous):

-90?

OpenStudy (anonymous):

Last line is -sin(75). Where did you get 90 from?

OpenStudy (anonymous):

i have no idea what -sin 35= ?

OpenStudy (anonymous):

It's meant to be -sin(75) , sorry! But sin (75) = sin(45 + 30) = sin(45)cos(30) + sin(30)cos(45) And sin(285) = sin(-75) = -sin(75) = -(sin(45)cos(30) + sin(30)cos(45)) Because sin(-a) = -sin(a)

OpenStudy (anonymous):

the answer is?

OpenStudy (anonymous):

hmm, INew, wouldn't it be easier if you just drew the graph and compute it from the graph? :) it'll be alot easier ^_^

OpenStudy (anonymous):

sin(45) = cos(45) = 1/sqrt(2) sin(30) = 1/2 cos(30) = sqrt(3)/2 \[\implies -\sin(75) = -\frac{1}{\sqrt{2}} \cdot \left(\frac{1}{2} + \frac{\sqrt{3}}{2}\right) = -\frac{1+ \sqrt{3}}{2\sqrt{2}} = -\frac{\sqrt{2} + \sqrt{6}}{4} = \sin(285)\]

OpenStudy (anonymous):

Easy to compute \[-\frac{\sqrt{6} + \sqrt{2}}{4}\] from a graph? You must be pretty badass with graphs.

OpenStudy (anonymous):

lol, yes by drawing the graph. You know turning the lines in circles we have 90 -180 360 -270 he'll easily figure out angle by then :)

OpenStudy (anonymous):

any *

OpenStudy (anonymous):

Just admit you were wrong, please, and I won't have to own you.

OpenStudy (anonymous):

lol, own who?

OpenStudy (anonymous):

the graph is actually easier, never realized it :p

OpenStudy (anonymous):

LOL, do you know how yo use it loco :)?

OpenStudy (anonymous):

The question specifically asks to use the addition formulae, and besides, no 'circle' will give you the angle 285.

OpenStudy (anonymous):

lol, yes it will :) , you've never used it right?

OpenStudy (anonymous):

kind of, ta showed it to us. still iffy on it but practicing.

OpenStudy (anonymous):

besides loco, the information INew has given you also is valuable, work with both ^_^

OpenStudy (anonymous):

thanks guys :)

OpenStudy (anonymous):

np ^_^

OpenStudy (anonymous):

but did you figure it out ? :)

OpenStudy (anonymous):

sstarica, please, draw me a circle which can give the exact value I calculated without, in essence, just using the addition formulae in diagram form.

OpenStudy (anonymous):

I can't draw it here, but you can twirl the line till you get the given angle, then you can figure out which angles to add. Those angles will be easy to compute since they're small, and in the end you can just add them up to find the sin of them ^_^

OpenStudy (anonymous):

stop fighting guys

OpenStudy (anonymous):

Oh, adding angles? Weird, just what I did. BTW drawing is harder than saying 285 = 360 - 75 => we need -sin(75)

OpenStudy (anonymous):

Simplify the trigonometric expression. cos[x]/sec[x] + sin[x]/csc[x]

OpenStudy (anonymous):

Sorry, I thought they were implying they could do the hard part in a circle, not the easy part :/

OpenStudy (anonymous):

I'm not fighting LOL, we're discussing loco ^_^

OpenStudy (anonymous):

1/sec = cos. 1/csc = sin. cos * 1/sec = .... etc

OpenStudy (anonymous):

how about we i mean u guys solve this problem

OpenStudy (anonymous):

you still didn't solve it? .-.

OpenStudy (anonymous):

no, i hate this stuff

OpenStudy (anonymous):

don't say that, you still didn't master it , but you will with practice and it'll be easy on you ^_^

OpenStudy (anonymous):

i've been practicing a lot, but still having trouble :'(

OpenStudy (anonymous):

\[\frac{\cos x}{\sec x} = \frac{\cos x}{\frac{1}{\cos x}} = \cos^2x\]

OpenStudy (anonymous):

it's alright, keep on practicing and you'll master it :)

OpenStudy (anonymous):

sstarica help?

OpenStudy (anonymous):

hold on :)

OpenStudy (anonymous):

newton it's wrong

OpenStudy (anonymous):

What's wrong? (Before answering, note that I am never wrong)

OpenStudy (anonymous):

sorry, i thought that was the answer

OpenStudy (anonymous):

we all make mistakes Inew :) even I, even the most intelligent, but noone said you're wrong, but your way is complicated.

OpenStudy (anonymous):

try to explain it in a simple way :)

OpenStudy (anonymous):

I was doing one half of it. See if you can apply the same logic to the next part - you might find it simlpifies to.. something nice.

OpenStudy (anonymous):

He meant I am wrong about the second problem he posted in here. I only did half of it

OpenStudy (anonymous):

Yes, my FIRST method was comlpicated, but my second was amazing.

OpenStudy (anonymous):

he posted another question?

OpenStudy (anonymous):

lol, good that you've made it easier ^_^

OpenStudy (anonymous):

Thankfully still no graphs required

OpenStudy (anonymous):

oh that one , LOL LOCO! you've posted it in a different thread =D and I was writing it down

OpenStudy (anonymous):

u can write it here

OpenStudy (anonymous):

nvm it, INew answered it though :)

OpenStudy (anonymous):

it's not right though

OpenStudy (anonymous):

Alright we have :\[\cos(x)/\sec(x) + \sin(x)/\csc(x)\] sec x = 1/cosx and csc x = 1/sin x so: \[[\cos(x)/1/\cos(x)] + [\sin(x)/1/\sin(x)]\] simplify and you'll get : \[= \cos^2(x) +\sin^2(x)\] \[= 1\] Correct me if I'm wrong please ^_^

OpenStudy (anonymous):

you are right :) i got the same answer

OpenStudy (anonymous):

No, you aren't wrong. Unless you count not using \frac{}{} for fractions as wrong (I personally do)

OpenStudy (anonymous):

well done ^_^

OpenStudy (anonymous):

lol, I didn't get you INew

OpenStudy (anonymous):

\[\frac{1}{2} \geq 1/2\]

OpenStudy (anonymous):

hmm, I can't find the frac() symbol lol >_<

OpenStudy (anonymous):

still don't know how to make that fraction line ^_^"

OpenStudy (anonymous):

last problem.... Simplify the trigonometric expression. cos[x]/ tan[x] + sex[x]

OpenStudy (anonymous):

frac{a}{b} gives\[\frac{a}{b}\]

OpenStudy (anonymous):

oh, thank you :)

OpenStudy (anonymous):

sex(x) - lolololololol

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

sec[x], sorry :P

OpenStudy (anonymous):

wish math was that fun

OpenStudy (anonymous):

I general, always write tan as sin/cos , sec as 1/cos etc.

OpenStudy (anonymous):

When trying to simplify things.

OpenStudy (anonymous):

It's alright, we have : \[=\cos^2(x)/sinx + 1/sinx\] \[= [\cos^2x + 1]/\sin(x)\] \[= \sin^2x/sinx = \sin(x)\] ^_^

OpenStudy (anonymous):

ORLY?

OpenStudy (anonymous):

maybe? :) correct me lol

OpenStudy (anonymous):

it's not correct

OpenStudy (anonymous):

\[\cos^2x + 1 \not= \sin^2x\]

OpenStudy (anonymous):

hmm, where is my mistake

OpenStudy (anonymous):

LOL! right >_< sorry

OpenStudy (anonymous):

For what it's worth, I don't think it simplifies much nicer than how it starts, so is a bad question.

OpenStudy (anonymous):

hmm, solve it :)

OpenStudy (anonymous):

man, u guys are at it.

OpenStudy (anonymous):

LOL, loco we're not fighting =P, hold on :)

OpenStudy (anonymous):

Ugh, it doesn't simplify.. at all. But also sec x = 1/cos(x) ... >_> sorry